OpenStudy (anonymous):

On Monday, the temperature was 50 degrees. On Tuesday, the temperature was 61 degrees. What is the percent increase?

OpenStudy (anonymous):

it is 81.9672131%

OpenStudy (texaschic101):

percent increase = (new amount - original amount) / original amount...then multiply that by 100 to get your percent. (61 - 50) / 50 = 11/50 = 0.22 0.22 * 100 = 22% increase

OpenStudy (anonymous):

22%